# NaturalReasoning: Reasoning in the Wild with 2.8M Challenging Questions

### v2

@@ Abstract
Scaling reasoning capabilities beyond traditional domains such as math and coding is hindered by the lack of diverse and high-quality questions. To overcome this limitation, we introduce a scalable approach for generating diverse and challenging reasoning questions, accompanied by reference answers. We present NaturalReasoning, a comprehensive dataset comprising 2.8 million questions that span multiple domains, including STEM fields (e.g., Physics, Computer Science), Economics, Social Sciences, and more. We demonstrate the utility of the questions in NaturalReasoning through knowledge distillation experiments which show that NaturalReasoning can effectively elicit and transfer reasoning capabilities from a strong teacher model. Furthermore, we demonstrate that NaturalReasoning is also effective for unsupervised self-training using external reward models or self-rewarding.
@@ 1 Introduction
Large language models (LLMs) have recently demonstrated increased reasoning capabilities to solve complex tasks. Notable examples include OpenAI’s o1 models (OpenAI et al., 2024) and DeepSeek’s R1 models (Guo et al., 2025), which have demonstrated significant improvements in challenging reasoning benchmarks such as competition math competitive coding, and GPQA (Rein et al., 2024). These models are designed to devote more time to deliberation before responding, enabling them to tackle intricate tasks and solve more complex problems in science, coding, and mathematics.
However, existing reasoning datasets often have limitations, primarily consisting of problem-solving tasks that exclude domains with open-ended reasoning and that lack diversity in scale and difficulty. To address this gap, we introduce NaturalReasoning, a comprehensive dataset curated from pretraining corpora, comprising 2.8 million reasoning questions spanning various topics, including Mathematics, Physics, Computer Science, Economics & Business, etc. NaturalReasoning is compared to a wide range of typical reasoning datasets, showcasing its advantageous properties, in particular its diversity and difficulty.
NaturalReasoning possesses several desirable attributes as a dataset, serving multiple research purposes. Firstly, the questions are backtranslated from the pretraining corpora, ensuring that it represents real-world reasoning problems that people care about, as opposed to synthetic datasets derived from benchmark datasets like MetaMathQA (Yu et al., 2024a) and OpenMathInstruct-2 (Toshniwal et al., 2024b). Secondly, it consists of questions with both verifiable and open-ended answers (e.g., theorem proving), providing a rich resource for developing learning algorithms to enhance LLMs’ reasoning abilities beyond a narrow subset of verifiable questions. Lastly, we demonstrate that the questions in NaturalReasoning require more advanced reasoning than existing datasets, and thus can be instrumental in improving LLMs’ reasoning capabilities through knowledge distillation from a stronger teacher model to a weaker student model or self-training based on external reward models or self-rewarding techniques (Yuan et al., 2024a).
@@
We create a large-scale and high-quality reasoning dataset by using pretraining data and LLMs alone without extra human annotation. The dataset contains challenging questions which require deliberate thinking accompanied with short reference answers.
We show that NaturalReasoning is a highly performant dataset to enhance reasoning capabilities in post-training. Specifically, using questions in NaturalReasoning in distillation is more sample efficient than existing datasets.
We explore self-training methods using the model’s own responses. Our results show that the questions in our dataset effectively enable self-learning, where self-rewarding techniques can even outperform strong external reward models.
@@ 2 Data Collection
| OpenMath2 | Math | 607K | 4644 | Synthesized by Llama3-405B-Inst from GSM8K+MATH | Llama3.1-405B-Instruct |
@@
| NaturalReasoning | Multi | 2.8M | 5521 | Synthesized by Llama3.3-70B-Inst grounded in Web | Llama3.3-70B-Instruct |
Backtranslating questions based on pretraining corpora has shown to be a scalable approach to create instruction-following datasets (Li et al., 2023; Yue et al., 2024). We take a similar approach to extract diverse and realistic reasoning questions grounded in pretraining data. An overview of the data creation approach is illustrated in Figure 1. A key differentiator of our approach is its emphasis on simplicity; we use LLMs throughout the entire cycle of synthetic data generation and curation, without any human annotation nor manual steps such as preprocessing to detect relevant documents or extracting questions with rule-based filtering.
2.1 Annotate Reasoning
Inspired by the meta-cognitive capabilities of state-of-the-art LLMs (Didolkar et al., 2024), we use an LLM to perform detailed annotation of pretraining documents to detect documents which contain sophisticated reasoning traces. We use the public pretraining corpora DCLM-baseline (Li et al., 2024c) and FineMath (Liu et al., 2024b) as sources, which have demonstrated steeper scaling laws than alternative corpora. More specifically, given a document from the pretraining corpus, we prompt an LLM to rate the content in along multiple axes: Problem Completeness, Problem Complexity and Technical Depth, Technical Correctness and Accuracy, Thinking and Reasoning. Empirically, we found that the model could analyze the document well and is able to follow the format requirement in the instruction. Please refer to Appendix D for the detailed prompt we use.
2.2 Synthesize Questions, Answers and Responses
For documents which are identified with a high degree of reasoning (e.g., scored full points on the axes of Problem Complexity and Technical Depth, Thinking and Reasoning), we further prompt an LLM to compose a self-contained and challenging reasoning question based on the content in . Different from existing work, which extracts questions appearing in the document (Yue et al., 2024), our approach allows us to synthesize more novel questions not directly contained in pretraining corpora. Additionally, we prompt the LLM to verify whether a correct reference answer to the synthesized question can be derived from and, if possible, include it as a reference answer. Besides questions and reference answers, we also augment each question in our dataset with a response generated by a relatively strong open-sourced model, Llama3.3-70B-Instruct.
@@ 2.3 Question Deduplication and Decontamination
Deduplication
Our deduplication process focuses on identifying and removing near-duplicate questions using locality-sensitive min-hashing at the word level. We apply a similarity threshold of 0.55 to filter out closely related variations, ensuring that questions with the same core reasoning task but different prompts are not redundantly included.
Decontamination
We filter out questions that are similar to popular reasoning benchmarks including MATH (Hendrycks et al., 2021b), GPQA, MMLU-Pro (Wang et al., 2024) and MMLU-STEM (Hendrycks et al., 2021a). The standard 13-gram decontamination method from EleutherAI’s llm-eval-harness (Gao et al., 2024) is used to identify and remove items from the dataset.
@@ 3 Data Analysis
MetaMathQA
is created by bootstrapping mathematical questions from GSM8K and MATH, rewriting them from multiple perspectives to form a new dataset (Yu et al., 2024b). The responses are generated using GPT-3.5-Turbo.
OpenMathInstruct-2
WebInstruct
recalls relevant documents from Common Crawl using a fastText model trained on a diverse seed dataset of quiz websites. It then extracts question-answer pairs contained in recalled web pages and uses LLMs (Qwen-72B (Bai et al., 2023), Mixtral-8×7B (Jiang et al., 2024)) to refine the answer (Yue et al., 2024).
@@ 3.1 Basic Statistics
We present a comparison of key dataset statistics in Table 1. Most large open reasoning datasets primarily focus on the math domain, with datasets such as OpenMathInstruct-2, and MetaMathQA containing only math-related questions. In contrast, NaturalReasoning covers reasoning problems from more diverse domains. Additionally, NaturalReasoning consists of 2.8M unique questions, significantly larger than OpenMathInstruct-2 (607K), and MetaMathQA (395K), though smaller than WebInstruct (13M).
Moreover, our questions have an average length of 55 words, which is longer than those in OpenMathInstruct-2 (46), WebInstruct (34), and MetaMathQA (41). The increased question length suggests that our dataset contains more complex and detailed problems, potentially requiring more advanced reasoning capabilities. This, combined with our diverse question sources and grounding in web-based knowledge, highlights the uniqueness and potential difficulty of our dataset.
@@ 3.2 Question Quality and Difficulty
To further evaluate questions in NaturalReasoning, we analyze them through the lens of both quality and difficulty. For each question in NaturalReasoning, we further annotate its quality (e.g., solvability, completeness) on a scale of 1 to 10 using Llama3.3-70B-Instruct. For the datasets we compare to, we randomly selected 10% of their questions and applied the same annotation prompt to assess their quality. The quality scores range from 0 to 10, where we categorize scores between 0 and 6 as low quality and those between 7 and 10 as high quality. The results, presented in Figure 2, show that NaturalReasoning has the highest proportion of high-quality questions at 98%, while the second-highest dataset contains 87% high-quality questions.
@@
To estimate question difficulty, we leverage a strong LLM to generate responses and use response length as a proxy, as longer chain-of-thoughts typically correspond to more complex questions. Specifically, we randomly selected 10% of questions from each dataset, and employ Llama3.3-70B-Instruct to generate responses for each question. As is shown in Figure 3, NaturalReasoning exhibits the longest median response length (434 words), significantly surpassing all other datasets. The second-longest median response length is observed in OpenMathInstruct-2, with a median of 354 words. This suggests that our dataset contains more intricate and reasoning-demanding questions compared to existing open reasoning datasets.
@@ 3.3 Question Diversity
We use an off-the-shell sentence encoder111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to generate embeddings for questions in WebInstruct and NaturalReasoning. We then apply UMAP (McInnes et al., 2018) to project the high-dimensional embeddings into a 2D space, followed by K-means clustering (Wu, 2012) to identify distinct groups. We use Mixtral-8B to assign high-level labels to these clusters, which is prompted with a few representative examples from each cluster. As is shown in Appendix LABEL:fig:topic_clusters, NaturalReasoning contains a more diverse and dense representation of non-mathematical topics, including Physics, Chemistry, Computer Science, and Law, besides Math. In contrast, WebInstruct is primarily skewed toward mathematical content, highlighting the broader topic coverage of NaturalReasoning.
@@
To estimate the topic distribution, a multi-class topic classifier is used to classify each question into 16 knowledge classes. The class labels are motivated by Wikipedia academic disciplines. Figure 4 shows that NaturalReasoning is complementary to WebInstruct, where NaturalReasoning has better coverage on non-Math topics especially Physics, Computer Science, Social Science, etc.
@@ 3.4 Reference Answer Analysis
Among the 2.8 million questions we synthesized, 81.68% have reference answers which could be derived from pretraining data. The distribution of reference answer lengths is illustrated in Figure 5, with single-word answers accounting for 10.7%, short answers (2–9 words) making up 20.0%, and long answers (10 words) constituting the majority at 50.9%.
We provide examples of questions with single-word, short, and long answers in Appendix B. In general, we found that questions with single word answers typically involve numerical, factual, or definitional queries, while questions with long answers demand more free-form in-depth analysis. For questions with a long answer, the extracted reference answer is typically a short summary content from the original documents or useful clues to answer the question.
While reference answers may contain some noise, we demonstrate their usefulness in section 6.
4 Experiment Setup
4.1 Training data
We highlight the efficacy of NaturalReasoning in two primary ways: (1) Knowledge distillation and supervised finetuning, which enables steeper scaling trends than other existing reasoning datasets, and (2) using NaturalReasoning as a source to extract seed data to target a specific domain at hand.
To demonstrate (2), we focus on science reasoning benchmarks, such as GPQA. To achieve this, we first sample 250 questions from the GPQA dataset, excluding those from the Diamond subset. For each selected question, we retrieve the 1,000 most similar questions from NaturalReasoning, which were already decontaminated against the entire GPQA datset. Similarity is computed using cosine similarity between two question embeddings. After obtaining this candidate set, we apply deduplication and perform clustering, grouping the questions into 15,000 clusters. From each cluster, we select the questions closest to the cluster center, ensuring a diverse and representative dataset for downstream science reasoning tasks. This process resulted in a pool of 15,000 questions, which we refer to as science-reasoning-15k. Models trained on this subset can still be evaluated on the GPQA-Diamond test set as those questions are not used for data selection. science-reasoning-15k is used in section 6 and section 7.
4.2 Evaluation
@@ 5 Steeper Scaling with Challenging and Diverse questions
Our hypothesis is that challenging and diverse questions which require thinking and reasoning are more sample efficient for post-training. To verify this, we run supervised finetuning (SFT) starting from a base model, and evaluate general reasoning capabilities and overall performance across MATH, GPQA, and MMLU-Pro.
We fine-tuned the Llama3.1-8B-Base model using fairseq2 training recipes (Balioglu, 2023), exploring the impact of varying dataset sizes (measured by the number of unique questions). Specifically, we trained the model on our dataset NaturalReasoning and the comparison datasets introduced in Table 1, and evaluated its average performance across three benchmarks: MATH, GPQA, and MMLU-Pro. For all datasets, we train 3 epochs for each data size, using batch size 128, learning rate , with cosine learning rate schedule where final learning rate is 1% of peak learning rate.

### v3

@@ Abstract
Scaling reasoning capabilities beyond traditional domains such as math and coding is hindered by the lack of diverse and high-quality questions. To overcome this limitation, we introduce a scalable approach for generating diverse and challenging reasoning questions, accompanied by reference answers. We present NaturalReasoning, a comprehensive dataset comprising 2.8 million questions that span multiple domains, including STEM fields (e.g., Physics, Computer Science), Economics, Social Sciences, and more. We demonstrate the utility of the questions in NaturalReasoning through knowledge distillation experiments which show that NaturalReasoning can effectively elicit and transfer reasoning capabilities from a strong teacher model. Furthermore, we demonstrate that NaturalReasoning is also effective for unsupervised self-training using external reward models or self-rewarding. To foster future work, we publicly release NaturalReasoning at https://huggingface.co/datasets/facebook/natural_reasoning.
@@ 1 Introduction
Large language models (LLMs) have demonstrated increased reasoning capabilities (OpenAI et al., 2024; Guo et al., 2025). These models are designed to devote more time to deliberation before responding, enabling them to tackle intricate tasks and solve more complex problems in science, coding, and mathematics. Such reasoning models are trained via large-scale reinforcement learning on tasks where the reward can be derived using rule-based verification. Existing reasoning datasets are often limited to narrow domains where the solutions are short and easy to verify, while the majority of reasoning problems across broader domains are open-ended reasoning. To bridge this gap, we introduce NaturalReasoning, a comprehensive dataset curated from pretraining corpora, comprising 2.8 million reasoning questions spanning various topics, including Mathematics, Physics, Computer Science, Economics & Business, etc. NaturalReasoning is compared to a wide range of reasoning datasets, showcasing its advantageous properties, in particular its diversity and difficulty.
NaturalReasoning possesses several desirable attributes as a dataset, serving multiple research purposes. Firstly, the questions are backtranslated from the pretraining corpora, ensuring that it represents diverse reasoning problems in the real world, as opposed to synthetic datasets derived from benchmark datasets like MetaMathQA (Yu et al., ) and OpenMathInstruct-2 (Toshniwal et al., ). Secondly, it consists of both questions with easy-to-verify answers and those with open-ended solutions (e.g., theorem proving), providing a rich resource for developing learning algorithms to enhance LLMs’ reasoning abilities across broader domains.Thirdly, we show that NaturalReasoning poses more difficult reasoning problems than existing datasets. Its questions therefore provide an effective testbed for improving LLM reasoning—whether through knowledge distillation from a stronger teacher model or reinforcement learning with external and self-generated reward signals (Yuan et al., ). Lastly, the NaturalReasoning dataset complements existing reasoning datasets in terms of both quality and quantity.
@@
We create a large-scale and high-quality reasoning dataset by using pretraining data and LLMs alone without extra human annotation. The dataset contains challenging questions which require deliberate thinking accompanied with reference answers. We release the NaturalReasoning dataset to the research community at https://huggingface.co/datasets/facebook/natural_reasoning.
We show that NaturalReasoning is a highly performant dataset to enhance reasoning capabilities in post-training. Specifically, using questions from NaturalReasoning in distillation is more sample efficient than existing datasets.
We investigate how NaturalReasoning supports self-training methods. Our results show that the questions in our dataset effectively enable self-learning, where self-rewarding techniques can yield performance comparable to some strong external reward models.
@@ 2 Data Collection
Backtranslating questions based on pretraining corpora has been shown to be a scalable approach to create instruction-following datasets (Li et al., 2023; Yue et al., 2024). We take a similar approach to extract diverse and realistic reasoning questions grounded in pretraining data, i.e. we generate grounded synthetic data. An overview of the data creation approach is illustrated in Figure 1. A key differentiator of our approach is its emphasis on simplicity; we use LLMs throughout the entire cycle of synthetic data generation and curation, without any human annotation nor manual steps such as preprocessing to detect relevant documents and extracting questions with rule-based filtering.
2.1 Annotate Reasoning
Inspired by the meta-cognitive capabilities of state-of-the-art LLMs (Didolkar et al., 2024), we use an LLM to perform detailed annotation of pretraining documents to detect documents which contain sophisticated reasoning traces. We use the public pretraining corpora DCLM-baseline (Li et al., 2024b) and FineMath (Allal et al., 2025) as sources, which have demonstrated steeper scaling laws than alternative corpora. More specifically, given a document from the pretraining corpus, we prompt an LLM to rate the content in along multiple axes: Problem Completeness, Problem Complexity and Technical Depth, Technical Correctness and Accuracy, Thinking and Reasoning. The detailed prompt is provided in Appendix G. Empirically, we found that the model could analyze the document well and is able to follow the format requirement in the instruction.
@@ 2.1 Annotate Reasoning
| NuminaMath | Math | 860K | 4832 | Grade-level and competition-level math datasets | GPT-4o |
| OpenMath2 | Math | 607K | 4644 | Synthesized by Llama3-405B-Inst from GSM8K+MATH | Llama3-405B-Inst |
@@
| NaturalReasoning | Multi | 2.8M | 5521 | Synthesized by Llama-3.3-70B-Inst grounded in Web | Llama-3.3-70B-Inst |
2.2 Synthesize Questions and Answers
For documents which are identified with a high degree of reasoning (e.g., scored full points on axes of Problem Complexity and Technical Depth, Thinking and Reasoning), we further prompt an LLM to compose a self-contained and challenging reasoning question based on the content in . Different from existing work, which extracts questions appearing in the document (Yue et al., 2024), our approach allows us to synthesize more novel questions not directly contained in pretraining corpora. Then, we prompt the LLM to verify whether a correct reference answer to the synthesized question can be derived from and, if possible, include it as a reference answer. Finally, for every question we generate an additional response with a relatively strong open-source model (Llama-3-70B-Instruct), which we later use as a teacher signal for knowledge distillation (see Section 5).
@@ 2.3 Question Deduplication and Decontamination
Deduplication Our deduplication process focuses on identifying and removing near-duplicate questions using locality-sensitive min-hashing at the word level. We apply a similarity threshold of 0.55 to filter out closely related variations, ensuring that questions with the same core reasoning task but different prompts are not redundantly included.
Decontamination We filter out questions that are similar to popular reasoning benchmarks including MATH (Hendrycks et al., 2021b), GPQA, MMLU-Pro (Wang et al., 2024) and MMLU-Stem (Hendrycks et al., 2021a). The standard 13-gram decontamination method from EleutherAI’s llm-eval-harness (Gao et al., 2024) is used to identify and remove items from the dataset.
@@ 3 Data Analysis
MetaMathQA is created by bootstrapping mathematical questions from GSM8K and MATH, rewriting them from multiple perspectives to form a new dataset (Yu et al., 2024). The responses are generated using GPT-3.5-Turbo.
NuminaMath is a comprehensive collection of 860K pairs of math problems and solutions (Li et al., 2024c). The questions cover multiple sources including grade-level questions and competition problems. The solutions in NuminaMath dataset are generated or rewritten by GPT-4o.
WebInstruct recalls relevant documents from Common Crawl using a fastText model trained on a diverse seed dataset of quiz websites. It then extracts question-answer pairs contained in recalled web pages and uses LLMs (Qwen-72B (Bai et al., 2023), Mixtral-8×7B (Jiang et al., 2024)) to refine the answer (Yue et al., 2024).
@@ 3.1 Basic Statistics
We present a comparison of key dataset statistics in Table 1. Most large open reasoning datasets primarily focus on the math domain, with datasets such as OpenMathInstruct-2, NuminaMath, and MetaMathQA containing only math-related questions. In contrast, NaturalReasoning covers reasoning problems from more diverse domains. Additionally, NaturalReasoning consists of 2.8M unique questions, significantly larger than OpenMathInstruct-2 (607K), NuminaMath (860K), and MetaMathQA (395K), though smaller than WebInstruct (13M).
With an average length of 55 words, questions in NaturalReasoning are notably longer than those in OpenMathInstruct-2 (46), WebInstruct (34), NuminaMath (48), and MetaMathQA (41). Longer questions embed richer context and multi-step requirements, demanding deeper reasoning. Coupled with our varied, web-grounded sources, this added complexity sets NaturalReasoning apart as a uniquely challenging dataset.
@@ 3.2 Question Quality and Difficulty
We evaluate question quality in NaturalReasoning with both automatic and human judgments. Three strong LLMs (DeepSeek-R1-Distill-Qwen-32B, Qwen2.5-72B-Instruct, Llama-3-70B-Instruct) independently score every question on a 0–10 scale reflecting solvability and completeness. For each comparison dataset, we randomly sample 10% of its questions and apply the same prompt. Scores of 0–6 are labeled low quality and 7–10 high quality; a question is deemed high quality when at least two models assign a high score. As shown in Figure 2, NaturalReasoning contains the highest fraction of high-quality questions (93%), surpassing the next-best dataset (79%). To corroborate these results, we conduct a small human study: two expert annotators independently score 100 randomly selected questions from each dataset, and we average their ratings. The pattern holds—NaturalReasoning achieves the top mean score (6.45) versus 5.92 for the second best. Full details are provided in Appendix D.
@@
To estimate question difficulty, we leverage a strong LLM to generate responses and use response length as a proxy, as longer chain-of-thoughts typically correspond to more complex questions. Specifically, we randomly selected 10% of questions from each dataset, and employ Llama3.3-70B-Instruct to generate responses for each question. As is shown in Figure 2, NaturalReasoning exhibits the longest median response length (434 words), significantly surpassing all other datasets. This suggests that our dataset contains more intricate and reasoning-demanding questions compared to existing open reasoning datasets.
@@ 3.3 Question Diversity
We use an off-the-shell sentence encoder111https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to generate embeddings for questions in WebInstruct and NaturalReasoning. We then apply UMAP (McInnes et al., 2018) to project the high-dimensional embeddings into a 2D space, followed by K-means clustering (Wu, 2012) to identify distinct groups. We use Mixtral-8B to assign high-level labels to these clusters, which is prompted with a few examples from each cluster. As Figure 5 shows, NaturalReasoning contains a more diverse and dense representation of non-mathematical topics, including Physics, Chemistry, Computer Science, and Law, besides Math. In contrast, WebInstruct is primarily skewed toward mathematical content, highlighting the broader topic coverage of NaturalReasoning.
@@
To estimate the topic distribution, a multi-class topic classifier is used to classify each question into 16 knowledge classes. The class labels are motivated by Wikipedia academic disciplines222https://en.wikipedia.org/wiki/Outline_of_academic_disciplines. 3(a) shows that NaturalReasoning is complementary to WebInstruct, where NaturalReasoning has better coverage on non-Math topics especially Physics, Computer Science, Social Science, etc.
@@ 3.4 Reference Answer Analysis
Among the 2.8 million questions we synthesized, 81.68% have reference answers which could be derived from pretraining data. The distribution of reference answer lengths is illustrated in 3(b), with single-word answers accounting for 10.7%, short answers (2–9 words) making up 20.0%, and long answers (10 words) constituting the majority at 50.9%.
We provide examples of questions with single-word, short, and long answers in Appendix B. In general, we found that questions with single word answers typically involve numerical, factual, or definitional queries, while questions with long answers demand more free-form in-depth analysis. For questions with a long answer, the extracted reference answer is typically a short summary content from the original documents or useful clues to answer the question. While reference answers may contain some noise, we demonstrate their potential usefulness in Appendix F.
4 Experimental Setup
We highlight the efficacy of NaturalReasoning in two settings: (1) Knowledge distillation, and (2) Unsupervised Self-Training. For (1), we evaluate whether NaturalReasoning enables steeper scaling than existing datasets when distilling reasoning capabilities to a student model via supervised finetuning (Section 5. We experiment with different model families such as Llama3.1-8B and Qwen2.5-7B. Specifically we show that questions from NaturalReasoning are very effective at distilling long chain-of-thoughts from reasoning models and compare it to manually curated questions such as LIMO (Ye et al., 2025) and S1K (Muennighoff et al., 2025) (Section 6). To demonstrate (2), we evaluate how well NaturalReasoning supports self-training either through a strong external reward model or self-rewarding mechanisms (Yuan et al., 2024) (Section 7).
Evaluation
@@ 5 Steeper Scaling with Challenging and Diverse questions
Our hypothesis is that challenging and diverse questions which require thinking and reasoning are more sample efficient for post-training. To verify this, we run supervised finetuning (SFT) starting from a base model, and evaluate overall performance across MATH, GPQA, and MMLU-Pro.
We fine-tuned the Llama3.1-8B-Base model and Qwen2.5-7B using fairseq2 training recipes (Balioglu, 2023), exploring the impact of varying dataset sizes. Specifically, we trained models on our dataset and the comparison datasets introduced in Table 1, and evaluated their average performance across three benchmarks: MATH, GPQA, and MMLU-Pro. For all datasets, we train 3 epochs for each data size, using batch size 128, learning rate , with cosine learning rate schedule where final learning rate is 1% of peak learning rate.